OpenStudy (anonymous):

Evaluate 3x2 − x, if x = -4. Answer (A) -52 (B) -44 (C) 44 (D) 52

ganeshie8 (ganeshie8):

put x = -4

ganeshie8 (ganeshie8):

3x^2 - x 3(-4)^2 - (-4)

ganeshie8 (ganeshie8):

3(16) + 4 48 + 4 = ?

OpenStudy (anonymous):

um so whats the Answer lol Im lost

ganeshie8 (ganeshie8):

hmm add 4 to 48 its tough to explain :(

myininaya (myininaya):

@ganeshie8 has done all of the work and has left one step for you @LAGIRL69 You can do 48+4

OpenStudy (vishweshshrimali5):

\(3x^2 - x = 3(-4)^2 - (-4) = 3 * 16 + 4 = ?\)

myininaya (myininaya):

If you have to just count from 4 units from 48

OpenStudy (anonymous):

@myininaya um 52?

myininaya (myininaya):

Right

OpenStudy (vishweshshrimali5):

@myininaya are u able to see my last answer in form of latex

OpenStudy (anonymous):

so @myininaya its -52 is the answer

OpenStudy (vishweshshrimali5):

i am asking this because I cannot........

OpenStudy (vishweshshrimali5):

@LAGIRL69 no its 52 = 48 + 4

OpenStudy (vishweshshrimali5):

correct answer should be D

myininaya (myininaya):

It is important to know order of operations on these @LAGIRL69 I said "right" to your first answer @vishweshshrimali5 no (and also I wanted her to say the final answer anyways :) )

OpenStudy (anonymous):

Ok so its D after i do all the Math thast what i find Its D?

OpenStudy (vishweshshrimali5):

yes \(after\) \(all\) \(the\) \(maths\) \(you\) \(do\)